### v2

@@ Abstract
Knowledge graph (KG) learning offers a powerful framework for generating new knowledge and making inferences. Training KG embedding can take a significantly long time, especially for larger datasets. Our analysis shows that the gradient computation of embedding is one of the dominant functions in the translation-based KG embedding training loop. We address this issue by replacing the core embedding computation with SpMM (Sparse-Dense Matrix Multiplication) kernels. This allows us to unify multiple scatter (and gather) operations as a single operation, reducing training time and memory usage. We create a general framework for training KG models using sparse kernels and implement four models, namely TransE, TransR, TransH, and TorusE. Our sparse implementations exhibit up to 5.3x speedup on the CPU and up to 4.2x speedup on the GPU with a significantly low GPU memory footprint. The speedups are consistent across large and small datasets for a given model. Our proposed sparse approach can be extended to accelerate other translation-based (such as TransC, TransM, etc.) and non-translational (such as DistMult, ComplEx, RotatE, etc.) models as well.
footnotetext: 1Department of Computer Science and Engineering, Texas A&M University, USA. Correspondence to: Md Saidul Hoque Anik <anik@tamu.edu>.
@@ 1 Introduction
We develop a comprehensive library based on our sparse formulation. This library consolidates most computations into several SpMM function calls, allowing optimized SpMM to directly accelerate the overall runtime of KGE training. We also discuss how to extend this concept to other non-translational models such as DistMult or ComplEx in Appendix C. We observe that SpTransX models significantly outperform established knowledge graph frameworks, such as TorchKGE and DGL-KE, particularly in terms of training time and GPU memory usage. For example, the average improvement in training time for the TransE model is illustrated in Table 1.
@@ 2 Background
Figure 2: Top three CPU intensive functions for various translation-based KGE models and datasets (indicated in bracket). The redness represents the popularity of a function among models. The dark red box indicates that the corresponding function is used in several different models. Blue/Purple indicates that the function is typically exclusive to the current model. The dark gray box indicates the dataset loading time. The light gray box indicates the rest of the training time.
@@ 3.3 Sparse Operations in Graph ML
Expressing graph operations through sparse linear algebra has proven highly effective for developing efficient and scalable graph learning algorithms. For example, the forward and backward propagation in graph convolutional networks (GCNs) and graph attention networks (GATs) can be optimized with sampled dense-dense matrix multiplication (SDDMM), sparse-dense matrix multiplication (SpMM), or their combination, known as FusedMM Fey & Lenssen (2019); Wang (2019); Rahman et al. (2021). Similarly, various graph embedding and visualization algorithms utilize SpMM and sparse-sparse matrix multiplication (SpGEMM). Consequently, popular graph machine learning libraries, such as PyTorch Geometric (PyG) Fey & Lenssen (2019) and DGL Wang (2019), rely on optimized implementations of SpMM, SpGEMM, and SDDMM available in vendor-provided libraries like cuSparse, MKL, or open-source libraries such as iSpLib Hoque Anik et al. (2024), FeatGraph Hu et al. (2020), and SparseTIR Ye et al. (2023). Despite the wide success of sparse operations in GNNs and graph embeddings, to the best of our knowledge, existing knowledge graph embedding libraries do not leverage sparse operations for training KGE models.
@@ 4.1 Sparse Approach
In the following subsections, we briefly discuss how to perform training in a sparse approach using SpMM instead of regular embedding extraction for several translation-based models. Both forward and backward propagation of our approach benefit from the efficiency of a high-performance SpMM (proof shown in Appendix G). This concept also extends broadly to various other knowledge graph embedding (KGE) methods as well, including DistMult, ComplEx, and RotateE (detailed formulations are provided in Appendix C). The sparsity of our formulation and related computational complexity are discussed in Appendix A and B.
@@ 4.7 SparseTransX Framework
We develop a general framework for SpTransX model training to enable efficient translation-based model training for large KG datasets. The framework is implemented using PyTorch 2.3 and consists of four modules, which are briefly described below.
@@ 4.7.2 Dataloaders
Our framework contains various dataloaders for shared and distributed training. It supports several standard knowledge graph formats, such as TTL, RDF, and CSV. Additionally, it contains a streaming dataset module for datasets that are too large to fit in the memory. When invoked, it creates an SQLite representation of the knowledge graph and stores the entity-index mapping in the database along with the triplets. All dataloaders connect to the sparse model input using a common interface.
4.7.3 Utilties and Core Functions
@@
4.7.4 Distributed Training
This module enables distributed training for SparseTransX using PyTorch DDP (Data Distributed Parallel) and FSDP (Fully Sharded Data Parallel) frameworks.
@@ 5 Experimental Setting
We implement the SpTransX models using PyTorch Framework and compare their total training time and GPU memory allocation with other well-known KG frameworks. We run these experiments for 7 datasets consisting of various sizes on a single CPU and a single GPU system separately. Several brief experiments demonstrating distributed training and scaling capacity of SpTransX framework is shown in Appendix E and F.
@@ 5.3 Training Loop
We prepare 11 separate scripts (SpTransE, SpTransR, SpTransH, SpTorusE, transe-torchkge, transr-torchkge, transh-torchkge, toruse-torchkge, transe-dglke, transr-dglke, and transe-pyg) to train the models on various datasets. Each script receives the dataset name as a command-line argument. The dataset is loaded from a shared repository. All frameworks use the same training configuration (learning rate: 0.0004, margin: 0.5), dissimilarity function (L2 or L2 torus), and loss function (MarginRankingLoss) and run for 200 epochs. Batch size and embedding dimensions are selected to maximize accuracy while utilizing available GPU memory (see subsection 6.1). The following table lists the dimensions and batch sizes used for different models.
Table 4: Training configuration for CPU and GPU
@@ 5.5 SparseTransX Configuration
In the SparseTransX framework, users can choose to use any high-performance SpMM to perform model training. We select iSpLib SpMM Hoque Anik et al. (2024) for CPU training and DGL g-SpMM Wang (2019) for GPU training.
@@ 6.1 Hyperparameter Selection
Figure 5: Hits@10 accuracy w.r.t. embedding size for FB15K dataset. 100 epoch training with batch size of 32768 and relation entity dimension as 8 (for TransH model)
Figure 6: Training time and GPU memory allocation w.r.t. Batch Size. 100 epoch training with entity dimension as 128 and relation dimension as 8 (for TransH model)
@@ 6.2.5 Model Accuracies
The sparse approach does not change the computational steps and thus does not affect the model accuracy. The accuracies of our implementations are consistent with that of other models, such as TorchKGE. For 100 epochs training on WN18 datasets with a fixed learning rate of 0.0004, SpTransX’s TransE, TorusE, and TransH models receive 0.72, 0.63, and 0.59 Hits@10 scores, whereas TorchKG’s models receive 0.74, 0.63, and 0.60. A more detailed evaluation (discussed in Appendix D) reveals that SpTransX achieves similar or better Hits@10 accuracy compared to TorchKGE when the training loop is equipped with a learning rate scheduler.
@@ 7 Conclusion
Despite the inherent sparsity of knowledge graphs and their embedding algorithms, existing frameworks often do not leverage sparse matrix operations to accelerate the training of KGE models. We have developed sparse formulations of translation-based KGE models that significantly outperform established knowledge graph frameworks, such as TorchKGE and DGL-KE, particularly regarding training time and GPU memory usage. Our findings demonstrate that the proposed approach consistently achieves improved performance across a range of both small and large datasets.
@@
Appendix A Applicability of SparseTransX for dense graphs
@@
Appendix B Computational Complexity
@@
Appendix C Applicability to Non-translational Models
@@
Figure 10: Loss curve for sparse and non-sparse approach. Sparse approach eventually reaches the same loss value with similar Hits@10 test accuracy.
Appendix D Model Performance Evaluation and Convergence
@@
Appendix E Distributed SpTransX and Its Applicability to Large KGs
SpTransX framework includes several features to support distributed KGE training across multi-CPU, multi-GPU, and multi-node setups. Additionally, it incorporates modules for model and dataset streaming to handle massive datasets efficiently.
Distributed SpTransX relies on PyTorch Distributed Data Parallel (DDP) and Fully Sharded Data Parallel (FSDP) support to distribute sparse computations across multiple GPUs.
Table 9: Average Time of 15 Epochs (seconds). Training time of TransE model with Freebase dataset (250M triplets, 77M entities. 74K relations, batch size 393K) on 32 NVIDIA A100 GPUs. FSDP enables model training with larger embedding when DDP fails.
Embedding Size DDP (Distributed Data Parallel) FSDP (Fully Sharded Data Parallel)
16 65.07 ± 1.641 63.35 ± 1.258
20 Out of Memory 96.44 ± 1.490
We run an experiment with a large-scale KG to showcase the performance of distributed SpTransX. Freebase (250M triplets, 77M entities. 74K relations, batch size 393K) dataset is trained using the TransE model on 32 NVIDIA A100 GPUs of NERSC using various distributed settings. SpTransX’s Streaming dataset module allows fetching only the necessary batch from the dataset and enables memory-efficient training. FSDP enables model training with larger embedding when DDP fails.
Appendix F Scaling and Communication Bottlenecks for Large KG Training
Communication can be a significant bottleneck in distributed KGE training when using SpMM. However, by leveraging Distributed Data-Parallel (DDP) in PyTorch, we successfully scale distributed SpTransX to 64 NVIDIA A100 GPUs with reasonable efficiency. The training time for the COVID-19 dataset with 60,820 entities, 62 relations, and 1,032,939 triplets is in Table 10.
Table 10: Scaling TransE model on COVID-19 dataset

### v3

@@ Abstract
Knowledge graph (KG) learning offers a powerful framework for generating new knowledge and making inferences. Training KG embedding can take a significantly long time, especially for larger datasets. Our analysis shows that the gradient computation of embedding is one of the dominant functions in the translation-based KG embedding training loop. We address this issue by replacing the core embedding computation with SpMM (Sparse-Dense Matrix Multiplication) kernels. This allows us to unify multiple scatter (and gather) operations as a single operation, reducing training time and memory usage. We create a general framework for training KG models using sparse kernels and implement four models, namely TransE, TransR, TransH, and TorusE. Our sparse implementations exhibit up to 5.3x speedup on the CPU and up to 4.2x speedup on the GPU with a significantly low GPU memory footprint. The speedups are consistent across large and small datasets for a given model. Our proposed sparse approach can be extended to accelerate other translation-based (such as TransC, TransM, etc.) and non-translational (such as DistMult, ComplEx, RotatE, etc.) models as well. An implementation of the SpTransX framework is publicly available as a Python package in https://github.com/HipGraph/SpTransX.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
We develop a comprehensive library based on our sparse formulation. This library consolidates most computations into several SpMM function calls, allowing optimized SpMM to directly accelerate the overall runtime of KGE training. We also discuss how to extend this concept to other non-translational models such as DistMult Yang et al. (2014a) or ComplEx Trouillon et al. (2016) in Appendix D. We observe that SpTransX models significantly outperform established knowledge graph frameworks, such as TorchKGE and DGL-KE, particularly in terms of training time and GPU memory usage. For example, the average improvement in training time for the TransE model is illustrated in Table 1.
@@ 2 Background
Figure 2: Top three CPU intensive functions for various translation-based KGE models and datasets (indicated in brackets). The redness represents the popularity of a function among models. The dark red box indicates that the corresponding function is used in several different models. Blue/Purple indicates that the function is typically exclusive to the current model. The dark gray box indicates the dataset loading time. The light gray box indicates the rest of the training time.
@@ 3.3 Sparse Operations in Graph ML
Expressing graph operations through sparse linear algebra has proven highly effective for developing efficient and scalable graph learning algorithms. For example, the forward and backward propagation in graph convolutional networks (GCNs) and graph attention networks (GATs) can be optimized with sampled dense-dense matrix multiplication (SDDMM), sparse-dense matrix multiplication (SpMM), or their combination, known as FusedMM Fey & Lenssen (2019); Wang (2019); Rahman et al. (2021). Similarly, SpMM and sparse-sparse matrix multiplication (SpGEMM) are widely used in algorithms for graph embedding Ranawaka et al. (2024); Rahman et al. (2020b), clustering Azad et al. (2018), and visualization Rahman et al. (2020a). Consequently, popular graph machine learning libraries, such as PyTorch Geometric (PyG) Fey & Lenssen (2019) and DGL Wang (2019), rely on optimized implementations of SpMM, SpGEMM, and SDDMM available in vendor-provided libraries like cuSparse, MKL, or open-source libraries such as iSpLib Hoque Anik et al. (2024), FeatGraph Hu et al. (2020), and SparseTIR Ye et al. (2023). Despite the wide success of sparse operations in GNNs and graph embeddings, to the best of our knowledge, existing knowledge graph embedding libraries do not leverage sparse operations for training KGE models.
@@ 4.1 Sparse Approach
In the following subsections, we briefly discuss how to perform training in a sparse approach using SpMM instead of regular embedding extraction for several translation-based models. Both forward and backward propagation of our approach benefit from the efficiency of a high-performance SpMM (proof shown in Appendix G). This concept also extends broadly to various other knowledge graph embedding (KGE) methods as well, including DistMult, ComplEx, and RotateE (detailed formulations are provided in Appendix D). The sparsity of our formulation and related computational complexity are discussed in Appendix B and C.
@@ 4.7 SparseTransX Framework
We develop a general framework for SpTransX model training to enable efficient translation-based model training for large KG datasets. The framework is implemented using PyTorch 2.3 and consists of several modules, which are briefly described below.
@@ 4.7.2 Dataloaders
Our framework contains various dataloaders for shared and distributed training. It supports several standard knowledge graph formats, such as TTL, RDF, and CSV. Additionally, it contains a streaming dataset module for datasets that are too large to fit in memory. When invoked, it creates an SQLite representation of the knowledge graph and stores the entity-index mapping in the database along with the triplets. All dataloaders connect to the sparse model input using a common interface.
4.7.3 Utilities and Core Functions
@@ 5 Experimental Setting
We implement the SpTransX models using PyTorch Framework and compare their total training time and GPU memory allocation with other well-known KG frameworks. We run these experiments for 7 datasets consisting of various sizes on a single CPU and a single GPU system separately. We provide a guideline to reproduce the experiment for one of the datasets in Appendix A. Although our framework supports distributed training using standard PyTorch data-parallel libraries, this paper focuses on sparse techniques and presents experiments conducted on a single GPU only. A preliminary result on distributed training of a large knowledge graph dataset, COVID-19 Tabassum et al. (2024), is shown in Appendix F.
@@ 5.3 Training Loop
We prepare 11 separate scripts (SpTransE, SpTransR, SpTransH, SpTorusE, transe-torchkge, transr-torchkge, transh-torchkge, toruse-torchkge, transe-dglke, transr-dglke, and transe-pyg) to train the models on various datasets. Each script receives the dataset name as a command-line argument. The dataset is loaded from a shared repository. All frameworks use the same training configuration (learning rate: 0.0004, margin: 0.5), dissimilarity function (L2 or L2 torus), and loss function (MarginRankingLoss), and run for 200 epochs. Batch size and embedding dimensions are selected to maximize accuracy while utilizing available GPU memory (see subsection 6.1). The following table lists the dimensions and batch sizes used for different models.
Table 4: Training configuration for CPU and GPU. Reduced embedding and batch size for TransR and TransH due to memory limitation.
@@ 5.5 SparseTransX Configuration
In the SparseTransX framework, users can choose to use any high-performance SpMM to perform model training. We select iSpLib SpMM Hoque Anik et al. (2024) for CPU training and DGL g-SpMM Wang (2019) for GPU training. CSR (Compressed Sparse Row) format is used for iSpLib, and COO (Coordinate) format is used for DGL as per the library requirement. The framework automatically generates sparse minibatches in the correct format when the user specifies the underneath SpMM library.
@@ 6.1 Hyperparameter Selection
Figure 5: Hits@10 accuracy w.r.t. embedding size for FB15K dataset. 100 epoch training with a batch size of 32768 and relation entity dimension as 8 (for TransH model). The TransH model encounters out-of-memory issues when the embedding size exceeds 256. Other models converge at an embedding size of approximately 2048 and show no improvement in Hits@10 accuracy for larger embeddings.
Figure 6: Training time and GPU memory allocation w.r.t. Batch Size. 100 epoch training with entity dimension as 128 and relation dimension as 8 (for TransH model).
@@ 6.2.5 Model Accuracies
The sparse approach does not change the computational steps and thus does not affect the model accuracy. The accuracies of our implementations are consistent with that of other models, such as TorchKGE. For 100 epochs training on WN18 datasets with a fixed learning rate of 0.0004, SpTransX’s TransE, TorusE, and TransH models receive 0.72, 0.63, and 0.59 Hits@10 scores, whereas TorchKG’s models receive 0.74, 0.63, and 0.60. A more detailed evaluation (discussed in Appendix E) reveals that SpTransX achieves similar or better Hits@10 accuracy compared to TorchKGE when the training loop is equipped with a learning rate scheduler.
@@ 7 Conclusion
Despite the inherent sparsity of knowledge graphs and their embedding algorithms, existing frameworks often do not leverage sparse matrix operations to accelerate the training of KGE models. We develop sparse formulations of translation-based KGE models that significantly outperform established knowledge graph frameworks, such as TorchKGE and DGL-KE, particularly regarding training time and GPU memory usage. Our findings demonstrate that the proposed approach consistently achieves improved performance across a range of both small and large datasets. We design a PyTorch-based library named SparseTransX that incorporates the sparse formulation methods for Knowledge Graph models and can demonstrate the aforementioned performance gains. The design of the library is flexible, and it can potentially perform distributed training once coupled with PyTorch Distributed Data Parallel (DDP) and Fully Sharded Data Parallel (FSDP) wrappers.
@@
Acknowledgements
This research is supported by the NSF OAC-2112606 and OAC-2339607 grants and DOE DE-SC0022098 and DE-SC0023349 awards.
Appendix A Artifact Evaluation
A.1 Abstract
We provide a guideline to reproduce the training times for one of the datasets reported in this paper. We share a publicly available GitHub repository (https://github.com/OnixHoque/sptransx-mlsys2025-reproduce) containing two bash scripts and a Jupyter Notebook that generate the workflow. The first bash script installs the frameworks needed to compare the performance with SpTransX and generates the environments required to run them. Another bash script runs the actual CPU and GPU experiments and stores the training times into text files. We include a validation Jupyter Notebook that performs processing on these text files and generates tables that depict Figure LABEL:fig:total_trg_time of the paper (for FB15K dataset). The experiments are run on one of the seven datasets (FB15K). The training time is computed for a single minibatch to keep the workflow short. The validation notebook approximates the total training time by multiplying the number of batches with the calculated training time. We further discuss how to run the workflow and perform validation from the generated results (also available in the Readme file of the GitHub repository). Please note that SpTransX is referred to as FastKG (former name) in the workflow.
A.2 Artifact check-list (meta-information)
  • Algorithm: TransE, TransR, TransH, TorusE (Knowledge Graph Embedding Training Models)
  • Program: PyTorch
  • Compilation: Conda, PIP, Python 3.7 and Python 3.8
  • Data set: FB15K Knowledge Graph Dataset
  • Run-time environment: Conda, GCC
  • Hardware: A CPU (preferably AMD) with 64 cores and 512GB DDR4 memory. An NVIDIA GPU with 40GB VRAM.
  • Run-time state: Training Time
  • Execution: Bash Script and Jupyter Notebook (requires CUDA for GPU experiments)
  • Metrics: Training Time (in seconds)
  • Output: Table showing total training time and slowdown factors of other frameworks
  • Experiments: Performs a single minibatch training on FB15K dataset for 200 epochs for all models available in all four frameworks in CPU and GPU
  • How much disk space required (approximately)?: 50 GB
  • How much time is needed to prepare workflow (approximately)?: 3 hours
  • How much time is needed to complete experiments (approximately)?: 3 hours
  • Publicly available?: Yes
  • Code licenses (if publicly available)?: MIT
  • Workflow framework used?: Bash Scripts and Jupyter Notebook
  • Archived (provide DOI)?: Not generated yet
A.3 Description
A.3.1 How delivered
The workflow is available in the following GitHub repository: https://github.com/OnixHoque/sptransx-mlsys2025-reproduce.
A.3.2 Hardware dependencies
The CPU and GPU experiments were run on dedicated CPU/GPU (single) nodes of NERSC Perlmutter. Their configurations are given below. The parameters are set to maximize CPU/GPU utilization. Similar configurations are recommended to reproduce the results. The CPU configuration is AMD EPYC 7763 (Milan) CPU with 64 cores and 512GB DDR4 memory. The GPU configuration is a single NVIDIA A100-SXM4 GPU with 40 GB VRAM.
A.3.3 Software dependencies
The experiments were tested on the following configuration.
  • GCC 12.2
  • Conda 24.9.1
  • Python 3.9 (3.8 for DGLKE)
  • PyTorch 2.3.1 (1.7.1 for DGLKE)
  • CUDAToolKit 12.1 (11.0 for DGLKE)
A.3.4 Datasets
FB15K dataset is used in the workflow. It is one of the seven datasets used in the paper. It is included in the repository.
A.4 Environment Installation
To set up the environments, clone the GitHub repository and run the following command. It will create two virtual environments. One specific to DGLKE, and another for the rest.
    ./0.setup_environments.sh
A.5 Experiment Workflow
To run the experiments, execute the following command.
    ./1.run_experiments.sh
It will generate the training time of a single minibatch training for various models and frameworks of the FB15K dataset. The outputs will be saved in cpu.txt and gpu.txt.
A.6 Evaluation and Expected Result
To generate the table of Figure LABEL:fig:total_trg_time (for FB15K) in the paper, execute the Jupyter Notebook 2.validation.ipynb. It will parse the generated text files and produce the tables for CPU and GPU for the FB15k dataset. The table includes the total training time and the slowdown factors of each framework compared to SpTransX.
It is expected that SpTransX is up to 5×5$\times$5 × and 4×4$\times$4 × faster in CPU and GPU, respectively, compared to other frameworks. The Jupyter Notebook contains the expected tables for both CPU and GPU from past experiments.
A.7 Notes
Please note that SpTransX is referred to as FastKG (former name) in the workflow.
Appendix B Applicability of SparseTransX for dense graphs
@@
Appendix C Computational Complexity
@@
Appendix D Applicability to Non-translational Models
@@
Figure 8: Loss curve for sparse and non-sparse approach. The sparse approach eventually reaches the same loss value with similar Hits@10 test accuracy.
@@
Appendix E Model Performance Evaluation and Convergence
Appendix F Preliminarily Experiment on Scaling
Communication can be a significant bottleneck in distributed KGE training when using sparse-dense matrix multiplication (SpMM) kernels. We perform a preliminary experiment to observe the scaling capacity of SparseTransX for the TransE model using a large knowledge graph dataset, COVID-19 Tabassum et al. (2024). We wrap the model in the Data Distributed Parallel (DDP) wrapper provided by PyTorch and train the model to 64 NVIDIA A100 GPUs to observe scaling behaviour. DDP (Distributed Data Parallel) trains a model by replicating it across multiple GPUs, synchronizing gradients during backpropagation to ensure consistent updates. The training times for various GPUs are reported in Table 9.
Table 9: Scaling TransE model on COVID-19 dataset with 60,820 entities, 62 relations, and 1,032,939 triplets Tabassum et al. (2024)